Suppose that in a certain neighborhood, the cost of a home (in thousands) is proportional to the size of the home in square feet. The regression equation quantifying this relationship is found to be (price) = 0.03*(size) + 35.06. You look more closely at one of the houses selected. The house is listed as having 2443 square feet and is listed at a price of $110.89 (thousand). What is the residual?

Question 4 options:

1)

2334.65

2)

2332.11

3)

-2332.11

4)

-2.54

5)

2.54
Answer #1

(price) = 0.03*(size) + 35.06

size = 2443

(price) = 0.03*(2443) + 35.06 = 108.35

residual = observed - predicted
= 110.89 - 108.35
= 2.54

Ans: Option 5